OpenStudy (anonymous):

Which of the following is the slope between the two points (4, 2) and (3, 1)? 1 -1 2 -2

OpenStudy (anonymous):

Y(2)-Y(1)/X(2)-X(1) that is your formula. Jut plug in the X's and Y's

hartnn (hartnn):

The slope of the line through points (x1,y1) and (x2,y2) is given by : \(\huge m=\frac{y_1-y_2}{x_1-x_2}\) now,just put the values and find m.

hero (hero):

insert the points into y = mx + b as follows: 1 = 3m + b 2 = 4m + b Isolate b in each equation b = 1 - 3m b = 2 - 4m set b = b 1 - 3m = 2 - 4m Place like terms on the same side 4m - 3m = 2 - 1 1m = 1 m = 1

OpenStudy (anonymous):

x would be 4 and 3/ and Y would be 2 and 1 right?

hartnn (hartnn):

yup

hero (hero):

If you use my method, you won't have to remember anything except y = mx + b

OpenStudy (anonymous):

3-1=2 4-2=2

hero (hero):

@Solmyr, you're confusing me bro. That's not how you do it.

OpenStudy (anonymous):

ok